OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (magbak):

the third equation

OpenStudy (magbak):

-2 -3 = -5

OpenStudy (mathstudent55):

First, find the slope of the line: m = (y2 - y1) / (x2 - x1) Then use the slope and one point to find the equation of the line using the equation below: y - y1 = m(x - x1)

OpenStudy (anonymous):

x - y = -5 i believe

OpenStudy (anonymous):

thank you all (:

OpenStudy (anonymous):

Your very welcome =)

OpenStudy (magbak):

You are welcome